# Model-Free Assessment of Simulator Fidelity via Quantile Curves

### v1

@@ Appendix A Technical Lemmas
We now choose
$\delta:=\frac{\eta}{2},\qquad c^{2}:=\frac{1}{2}\log\frac{2m}{\eta},$
so that
$\mathbb{P}(\mathcal{E}_{\mathrm{DKW}})\geq 1-\frac{\eta}{2},\qquad e^{-2c^{2}}=\frac{\eta}{2m}\Longrightarrow me^{-2c^{2}}=\frac{\eta}{2},$
and therefore
$\mathbb{P}\Big(\mathcal{E}_{\mathrm{DKW}}\cap\bigcap_{r=1}^{m}\mathcal{E}_{r}^{-}\Big)\geq 1-\eta.$
On this high-probability event, plugging $\alpha=\alpha_{r}$ and the choices of $(c,\delta)$ into (B.3) gives, for each $r$,
$\mathbb{P}_{\psi\sim\Psi}\big($ $\Delta(\psi)\geq\hat{V}_{m}^{-}(\tfrac{\alpha_{r}}{2})\big|\mathcal{D}\big)\geq$
$1-\alpha_{r}-\frac{c}{\sqrt{m}}\sqrt{4\alpha_{r}+\tfrac{4c^{2}+8}{m}}-\frac{2c^{2}+2}{m}-\varepsilon_{m}(\delta).$
A direct algebraic simplification, using $c^{2}=\tfrac{1}{2}\log(2m/\eta)$ and $\delta=\eta/2$, shows that the right-hand side is:
$1-\alpha_{r}-\frac{1}{\sqrt{m}}\varepsilon_{-}(\alpha,m,\eta),$
where
$\varepsilon_{-}(\alpha,m,\eta):=$ $\sqrt{2\alpha\log\frac{2m}{\eta}+\frac{\big(\log\frac{2m}{\eta}\big)^{2}+4\log\frac{2m}{\eta}}{m}}$
$+\frac{\log\frac{2m}{\eta}+2}{\sqrt{m}}+\sqrt{\frac{\log(4/\eta)}{2}},$
so that for each grid point $\alpha_{r}$ we can write
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta(\psi)\geq\hat{V}_{m}^{-}(\tfrac{\alpha_{r}}{2})\big|\mathcal{D}\big)\geq 1-\alpha_{r}-\frac{\varepsilon_{-}(\alpha_{r},m,\eta)}{\sqrt{m}}.$
Finally, for an arbitrary $\alpha\in(0,1)$, let $r:=\lceil m\alpha\rceil$ and $\alpha_{+}:=\alpha_{r}=r/m\in[\alpha,\alpha+1/m]$. Since $\hat{V}_{m}^{-}(\cdot)$ is nondecreasing,
$\hat{V}_{m}^{-}\big(\tfrac{\alpha}{2}\big)\leq\hat{V}_{m}^{-}\big(\tfrac{\alpha_{+}}{2}\big),$
andsince $\alpha_{+}\leq\alpha+1/m$ and $\varepsilon_{-}(\alpha_{+},m,\eta)$ differs from $\varepsilon_{-}(\alpha,m,\eta)$ by at most an $O(1)$ amount, we can enlarge $\varepsilon_{-}$ slightly (absorbing the $1/m$ rounding slack) so that
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\geq\hat{V}_{m}^{-}(\tfrac{\alpha}{2})\Big|\mathcal{D}\Big)\geq 1-\alpha-\frac{\varepsilon_{-}(\alpha,m,\eta)}{\sqrt{m}},$
for all $\alpha\in(0,1)$, on an event of probability at least $1-\eta$. This is exactly (5.1).
Finally, we derive the confidence band guarantee. Let $V$ denote the (left-continuous) quantile function of $\Delta(\psi)$,
$V(u):=\inf\{t\in\mathbb{R}:F_{\Delta}(t)\geq u\},$
where $F_{\Delta}$ is the CDF of $\Delta(\psi)$ under $\psi\sim\Psi$.
From Theorem 3.1, we already know that on the same high-probability event,
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\leq\hat{V}_{m}\big(1-\tfrac{\alpha}{2}\big)\Big|\mathcal{D}\Big)\geq 1-\alpha-\frac{\varepsilon(\alpha,m,\eta)}{\sqrt{m}},$ (B.12)
for some remainder $\varepsilon(\alpha,m,\eta)$ of the similar algebraic form as $\varepsilon_{-}(\alpha,m,\eta)$. Rewriting (5.1) and (B.12) in terms of $F_{\Delta}$ yields
$F_{\Delta}\Big(\hat{V}_{m}^{-}\big(\tfrac{\alpha}{2}\big)\Big)\leq\alpha+\frac{\varepsilon_{-}(\alpha,m,\eta)}{\sqrt{m}},$
$F_{\Delta}\Big(\hat{V}_{m}\big(1-\tfrac{\alpha}{2}\big)\Big)\geq 1-\alpha-\frac{\varepsilon_{(}\alpha,m,\eta)}{\sqrt{m}}$
By the definition of $V$ as the left-continuous inverse of $F_{\Delta}$, these inequalities are equivalent to
$\hat{V}_{m}^{-}\Big(\tfrac{\alpha}{2}\Big)\leq V\Big(\alpha+\tfrac{\varepsilon_{-}(\alpha,m,\eta)}{\sqrt{m}}\Big)$
$V\Big(1-\alpha-\tfrac{\varepsilon(\alpha,m,\eta)}{\sqrt{m}}\Big)\leq\hat{V}_{m}\Big(1-\tfrac{\alpha}{2}\Big).$
Combining these displays yields the desired quantile sandwich (5.2) for all $\alpha\in(0,1)$, on an event of probability at least $1-\eta$. This completes the proof.
@@ C.3 Synthetic Profile Generation
Education level ISCED levels 0–8333ISCED 0 corresponds to early childhood or no education; ISCED 8 corresponds to doctoral or equivalent.
@@ E.1 EEDI Dataset
We apply our methodology to produce a fidelity profile for each candidate LLM $\ell$. We use absolute error as the loss, L(p,q)=|pq|L(p,q)=|p-q|. We set $\gamma=0.5$ uniformly and the DKW failure probability to $\delta=0.1$, which determines the curve’s effective width at $\alpha\to 1$ in Figure 10. In addition, we set the simulation budget $k=50$.
Figure 10: Quantile fidelity profiles $\hat{V}(\alpha)$ across LLMs (Discrepancy: Absolute loss, $k=50$, $\beta=0.5$, $\delta=0.1$).
@@ E.2 OpinionQA
We apply our methodology to produce a fidelity profile for each candidate LLM $\ell$. We use total variation as the discrepancy measure, $L(p,q)=\tfrac{1}{2}\,\|p-q\|_{1}$, and set $\delta=0.05$. In addition, we set the simulation budget $k=100$, the result is presented in Figure 11.
@@
Figure 11 compares models by how tightly their synthetic outcomes track the human distribution across items. We plot $\hat{V}_{\ell}(\alpha)$ against $\alpha$, where lower-flatter curves indicate uniformly small discrepancies, while elbows reveal rare but severe misses. GPT-4o lies lowest across most quantiles, indicating the most reliable alignment, with Mistral 7B clearly performing worse. Notably, the simulator curves are steeper than the random benchmark, indicating question-dependent alignment and less uniform discrepancies. This suggests the simulators may require further fine-tuning to achieve more uniform discrepancy levels across this set of questions.

### v2

@@ B.3 Corollary for simulator budget:
For an arbitrary $\alpha\in(0,1)$, let $r=\lceil m\alpha\rceil$ and denote $\alpha_{+}:=\alpha_{r}=r/m\in[\alpha,\alpha+1/m]$. Applying (B.18) at $\alpha_{+}$ gives
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\geq\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha_{+},c,m)\big)\Big|\mathcal{D}\Big)\geq 1-\alpha_{+}-\varepsilon_{m}(\delta).$
Since $\alpha_{+}\geq\alpha$, we have $\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha_{+},c,m)\geq\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)$, and because $\hat{V}_{m}^{-}(\cdot)$ is nondecreasing,
$\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\leq\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha_{+},c,m)\big).$
Therefore,
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\geq\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\Big|\mathcal{D}\Big)\geq 1-\alpha_{+}-\varepsilon_{m}(\delta),$
and one may absorb the rounding slack $\alpha_{+}-\alpha\leq 1/m$ into the $O(m^{-1})$ term (as in Theorem 3.1). Hence, with probability at least $1-\delta-me^{-2c^{2}}$ over $\mathcal{D}$, the bound (B.18) holds uniformly for all $\alpha\in(0,1)$ (up to the $1/m$ discretization slack). This completes the proof.
@@ C.3 Synthetic Profile Generation
Education level ISCED levels 0–8444ISCED 0 corresponds to early childhood or no education; ISCED 8 corresponds to doctoral or equivalent.
@@
C.4 Adaptive coverage rate $\gamma_{j}$ choice
As mentioned in the main text, we focus on the case where we set the coverage rate to the generic functional form
$\gamma_{j}=1-n_{j}^{-\beta}.$
We consider a set of $\beta$ values in our sensitivity analysis. We conduct a numerical experiment over $\beta\in\{\tfrac{1}{5},\tfrac{1}{4},\tfrac{1}{3},\tfrac{1}{2}\}$ in Figure 12. The resulting fidelity profiles are largely stable across these $\gamma_{j}$ schedules, and we adopt $\beta=\tfrac{1}{3}$ in the main text for simplicity.
Figure 12: $\beta$ sensitivtiy analysis aunder $n=\{50,200\}$.
@@ E.1 EEDI Dataset
We apply our methodology to produce a fidelity profile for each candidate LLM. We use absolute error as the loss, L(p,q)=|pq|L(p,q)=|p-q| and set the simulation budget $k=50$.
Figure 13: Quantile fidelity profiles $\hat{V}(\alpha)$ across LLMs (Discrepancy: Absolute loss, $k=50$.
@@ E.2 OpinionQA
We apply our methodology to produce a fidelity profile for each candidate LLM. We use total variation as the discrepancy measure, $L(p,q)=\tfrac{1}{2}\,\|p-q\|_{1}$. In addition, we set the simulation budget $k=100$, the result is presented in Figure 14.
@@
Figure 14 compares models by how tightly their synthetic outcomes track the human distribution across items. We plot $\hat{V}_{\ell}(\alpha)$ against $\alpha$, where lower-flatter curves indicate uniformly small discrepancies, while elbows reveal rare but severe misses. GPT-4o lies lowest across most quantiles, indicating the most reliable alignment. Notably, the simulator curves are steeper than the random benchmark, indicating question-dependent alignment and less uniform discrepancies. This suggests the simulators may require further fine-tuning to achieve more uniform discrepancy levels across this set of questions. Nonetheless, the simulators are, in general, lower than random benchmark, which indicates we have a alignment result than EEDI.